OpenStudy (anonymous):

Just need an Ans. check please a) Find the eqation of the locus of a point which moves in the plane so that its distance from (-2,0) is always 3 times its distance from the origin. my Ans: x^2+y^2+4x-32=0 b) Show that the locus is a circle and state the center and radius of the circle. My Ans.: (x+2)^2+(y-0)^2=36 r^2=36=6 center(-2,0)

OpenStudy (anonymous):

distance fron the point is always 6 - the curve is a circle radius 6 centre -2,0 ( x + 2)^2 + y^2 = 36 x^2 + y^2 + 4x + 4 = 36 x^2 + y^2 + 4x + -32 = 0 you are correct

OpenStudy (anonymous):

sweet thnx

OpenStudy (anonymous):

since the curve is always the same ditsance from (-2,0) and this distance is 6, 6 is the radius of the circle and applying pythagoras gives the first line in the last post

OpenStudy (anonymous):

welcome